# Test-Time Anchoring for Discrete Diffusion Posterior Sampling

### v1

@@ 3.2.2 Anchored Remasking
Figure 2: Qualitative results on FFHQ and ImageNet for SR ($4\times$) and Gaussian deblur. Compared to DPS and G2D2, APS yields better results with sharper texture and refined facial features. For instance, in the third row, APS reconstructs fine strands of the white and brown dog’s fur.
@@
4 Experiments
Baselines. Since our focus is on discrete diffusion, we first compare against existing discrete posterior samplers: G2D2 (Murata et al., 2024) and SGDD (Chu et al., 2025)2). To provide a comprehensive evaluation, we also include established continuous baselines, both in pixel space (DPS (Chung et al., 2023), DDRM (Kawar et al., 2022), and DiffPIR (Zhu et al., 2023)) and in latent space (PSLD (Rout et al., 2023) and ReSample (Song et al., 2024)). This ensures our evaluation spans both discrete and continuous posterior sampling paradigms.
Benchmarks. We evaluate on the standard inverse problem benchmarks used in prior works. For high-resolution faces we use FFHQ at $256\times 256$ (Karras et al., 2019), and for diverse natural images we use ImageNet at $256\times 256$ (Deng et al., 2009). Performance is measured using three standard metrics: Learned Perceptual Image Patch Similarity (Zhang et al., 2018) (LPIPS), Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM) (Wang et al., 2004). All methods are evaluated at the same resolution and on the same images following G2D2 (Murata et al., 2024).
[…]
To demonstrate scalability, we upsample the benchmark images to $1024\times 1024$ and apply our posterior sampler to these scaled images (termed APS-L in discussion below). We defer implementation details to §B.1 and computational complexity to §B.5.
Table 2: Quantitative results on general inverse problems. We report results on two additional linear (random inpainting, motion deblurring) and nonlinear (HDR, nonlinear blur) tasks. Since G2D2 and SGDD do not evaluate on these tasks, we compare our discrete sampler against representative continuous baselines: DPS (pixel-space) and PSLD (latent-space).
@@ 4 Experiments
Figure 4: Qualitative results on stylization. We present four style–content combinations. For each case, our APS algorithm conditions on a single reference style image together with a text prompt to generate the stylized output images. The prompt follows the template: “Generate an image in [style] style. A [class], high detail, photorealistic.” Here, [style] denotes the reference style (e.g., Celestial Artwork), and [class] corresponds to the label shown below (e.g., Carousel).
@@ 4.1 Results on Linear Inverse Problems
Evaluation on FFHQ: Table 1 (a) shows that our approach outperforms prior discrete diffusion samplers across both super resolution and Gaussian deblurring tasks. For super resolution, APS reduces LPIPS by $13.65\%$ and improves PSNR by $2.11\%$. On Gaussian deblurring, APS lowers LPIPS by $3.83\%$ compared to G2D2 and raises PSNR by $5.88\%$. The large variant, APS-L, pushes performance even further, achieving up to a $31.36\%$ gain over G2D2 in terms of LPIPS. These results show that our quantized expectation (§3.2.1) and anchored remasking (§3.2.2) strategies not only outperform discrete diffusion baselines but often surpass strong continuous-diffusion methods, such as DiffPIR (which uses pixel-space diffusion) and PSLD (which uses latent-space diffusion), resulting in higher perceptual and reconstruction quality.
Figure 2 (top two rows) presents qualitative comparisons for super resolution ($4\times$) and Gaussian deblurring on FFHQ. DPS produces over-smoothed results with blurry facial details, while G2D2 often introduces artifacts and fails to restore a natural facial structure. Our APS sampler yields sharper textures and more faithful reconstructions, recovering details such as hair strands, facial contours, and eyeglass edges with higher perceptual quality. The large variant, APS-L, further enhances structure and realism, generating photo-realistic outputs with finer details and fewer artifacts. These results confirm that APS and APS-L deliver superior qualitative performance on face datasets, where perceptual fidelity is especially critical.
Evaluation on ImageNet: Table 1 (b) quantifies that our method achieves consistent improvements across both SR and Gaussian deblurring tasks on the ImageNet benchmark. For super resolution, APS reduces LPIPS by $7.16\%$ compared to G2D2 and improves PSNR by $4.74\%$. On Gaussian deblurring, APS improves PSNR by $8.81\%$ while maintaining comparable LPIPS. APS-L improves LPIPS by up to $35.82\%$ compared to G2D2. These results confirm that APS outperforms prior discrete posterior samplers and often surpasses continuous baselines such as DiffPIR and PSLD, delivering superior perceptual quality and reconstruction fidelity.
Figure 2 (bottom two rows) shows SR ($4\times$) and Gaussian deblurring results on ImageNet. Notably, DPS produces overly smooth outputs with a loss of fine details, while G2D2 introduces struggles to recover sharp edges. In contrast, APS reconstructs sharper textures (e.g., the fur of the dog and the feathers of the swan) and yields more natural color. The large variant, APS-L, further enhances structural fidelity, recovering finer details in challenging regions such as a frog’s skin texture and a goat’s fur. These examples highlight that our approach achieves superior perceptual quality and faithful structure reconstruction compared to both pixel-based and discrete diffusion baselines.
@@ 4.2 Results on General Inverse Problems
Table 2 shows that APS generalizes effectively to more challenging linear (random inpainting, motion deblurring) and nonlinear (HDR, nonlinear blur) inverse problems on FFHQ and ImageNet. Unlike prior discrete samplers such as G2D2 and SGDD, which were demonstrated only on limited tasks, APS consistently achieves strong perceptual quality (lower LPIPS) and reconstruction fidelity (higher PSNR). For instance, on ImageNet motion deblurring APS-L attains $0.318$ LPIPS and $25.19$ PSNR, substantially outperforming continuous baselines DPS and PSLD. Similarly, in nonlinear tasks such as HDR and nonlinear blur, APS delivers sharper, more consistent reconstructions, closing the gap with continuous diffusion while operating within a purely discrete framework. These results highlight the broader applicability and robustness of our approach compared to existing discrete diffusion samplers. Figure 3 shows the qualitative results of general inverse problems.
@@ 4.3 Results on Reference-based Stylization
Table 3: Quantitative results on stylization.
@@
We compare APS with the discrete diffusion baseline MMaDA (Yang et al., 2025) and continuous methods (shaded gray). Table 3 reports quantitative results, while Figure 4 shows qualitative examples. APS improves over MMaDA on ImageReward (Xu et al., 2024), CLIP-T (Radford et al., 2021), and DINO (Caron et al., 2021) scores, demonstrating that our posterior sampler unlocks stylization capabilities absent in the base model. Notably, APS also surpasses continuous baselines such as IP-Adapter (Ye et al., 2023) and StyleAligned (Hertz et al., 2023) on ImageReward and CLIP-T, despite relying on a weaker generative prior (Yang et al., 2025).
4.3.1 Results on Text-Guided Block Inpainting
We compare our approach against large-scale continuous diffusion baselines: Imagen3 (Baldridge et al., 2024), Flux (Black Forest Labs, 2024), and HDPainter (Manukyan et al., 2025). As illustrated in Figure 5, our method generates realistic clothing textures, with stronger alignment to the reference prompts and fewer artifacts (red boxes highlight failure regions of competing methods). In contrast, Imagen3 and Flux often introduce distorted or inconsistent garment regions, while HD-painter produces less faithful completions with mismatched styles.
[…]
The appendix provides extensive details on our Anchored Posterior Sampling method. We present the theoretical derivation of variational bounds (§A), all implementation specifics and hyperparameters (Algorithm 1, Table 4), and a detailed ablation study (§B.3, Figure 6) demonstrating the impact of our innovations. We also showcase the superior computational efficiency (Table 6) and numerous additional qualitative results (§B) across complex inverse problems, details of text-guided block inpainting (§B.8.2), and extended stylization results (Figure 12 and  13).

### v2

@@ 3 Test-Time Anchored Posterior Sampling
Figure 2: Qualitative results on FFHQ and ImageNet for SR ($4\times$) and Gaussian deblur. Compared to DPS and G2D2, APS yields better results with sharper texture and refined facial features. For instance, in the third row, APS reconstructs fine strands of the dog’s fur.
3.1 Quantized Expectation
Following the standard practice in test-time optimization (Chung et al., 2023; Rout et al., 2023, 2024; Chung et al., 2024; Rout et al., 2025c; Zhang et al., 2025; Murata et al., 2024), the upper bound in Theorem 3.3 is optimized at every time step $t$. We ignore the normalizing factor during optimization as discussed in Implication 3.2 and re-normalize the optimized logits via Softmax to construct a valid distribution. Since ${\mathcal{L}}_{\mathrm{NELBO}}({\mathbf{x}};\theta)$ does not depend upon $\varphi$, we obtain the following objective per time step:
${\mathcal{L}}_{t}(\varphi)$ $\coloneqq\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}\,\sum_{l=1}^{L}\Big[\log\frac{\langle{\mathbf{x}}^{l}_{\theta}({\mathbf{z}}_{t}),{\mathbf{x}}^{l}\rangle}{\langle{\mathbf{x}}^{l}_{\varphi}({\mathbf{z}}_{t}),{\mathbf{x}}^{l}\rangle}\,\bm{1}_{\{{\mathbf{z}}^{l}_{t}={\mathbf{m}}\}}\Big]$
$-\mathbb{E}_{q(Z_{s}|{\mathbf{z}}_{t},{\mathbf{x}})}\Big[\sum_{l=1}^{L}\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t};Z_{s}^{l}))\Big].$ (7)
Equation (3.1) remains intractable because the first term depends on ${\mathbf{x}}^{l}$ which is not known in posterior sampling and the second term requires ${\mathcal{O}}(K^{L})$ compute [^1] Recall that $\log q({\mathbf{y}}|{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t};Z_{s}^{l}))$ within the outer expectation of (3.1) is defined by an inner expectation in (6). The marginalization over $Z_{s}$ in the outer expectation costs ${\mathcal{O}}(2^{L})$ (given ${\mathbf{z}}_{t}$ and $\mathbf{x}$}, $Z_{s}\in\{{\mathbf{x}},{\mathbf{m}}\}$). For each $Z_{s}$, the computation of the inner expectation (over random variables except $Z_{s}^{l}$) inside the summation costs ${\mathcal{O}}(LK^{L-1})$. In the worst case, the total cost becomes ${\mathcal{O}}(K^{L})$.. We address the first term by replacing ${\mathbf{x}}^{l}$ with ${\mathbf{x}}^{l}_{\theta}({\mathbf{z}}_{t})$ because the pretrained model provides a reasonable approximation. For the second, since $q({\mathbf{y}}|{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t};Z_{s}^{l}))$ involves an expectation over a parameterized distribution, one could in principle, optimize ${\mathcal{L}}_{t}(\varphi)$ using policy-gradient methods such as REINFORCE (Williams, 1992), PPO (Schulman et al., 2017), or GRPO (Shao et al., 2024), as adopted in SVDD (Li et al., 2024). This leads to high-variance gradients and sparse rewards, resulting in poor sample quality (see: Table 1).
We introduce quantized expectation to better approximate the second term in (3.1). Recall that the measurement likelihood is modeled as $q({\mathbf{y}}|{\mathbf{x}})\propto\exp(-\tfrac{\|{\mathbf{y}}-{\mathcal{A}}({\mathcal{D}}({\mathbf{x}}))\|_{2}^{2}}{2\sigma^{2}}).$ Sampling ${\mathbf{x}}\sim{\mathbf{x}}_{\varphi}({\mathbf{z}}_{t})$ to compute likelihood makes ${\mathcal{L}}_{t}(\varphi)$ non-differentiable and noisy. Instead, we propose a differentiable and practically implementable surrogate [^2] Note that $\alpha_{t}-\alpha_{s}<0$; minimizing (3.1) is equivalent to maximizing the correlation between the prior and the tilted posterior.:
$\hat{{\mathcal{L}}}_{t}(\varphi)\coloneqq\frac{\alpha_{t}-\alpha_{s}}{1-\alpha_{t}}\,\sum_{l=1}^{L}$ $\Big[\log\langle{\mathbf{x}}^{l}_{\varphi}({\mathbf{z}}_{t}),{\mathbf{x}}^{l}_{\theta}({\mathbf{z}}_{t})\rangle\,\bm{1}_{\{{\mathbf{z}}^{l}_{t}={\mathbf{m}}\}}\Big]$
$-L\cdot\log q({\mathbf{y}}|\tilde{{\mathbf{x}}}_{\varphi}({\mathbf{z}}_{t}))\big].$ (8)
Appendix C.5.1 includes the derivation of (3.1) from (3.1). Differentiability has propelled posterior sampling to achieve state-of-the-art results using continuous diffusion (Chung et al., 2023; Rout et al., 2023, 2024; Zhang et al., 2025). To restore differentiability in discrete diffusion, we parameterize ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t(i)})=\mathrm{Softmax}(\varphi_{t(i)})$ with $\varphi_{t(i)}=\{\varphi^{l}_{t(i)}\}_{l=1}^{L}\in\mathbb{R}^{K\times L}$ containing logits $\varphi^{l}_{t(i)}$ over the codebook $\{{\mathbf{c}}_{k}\in{\mathcal{C}}\}$ for each position $l$. We compute the expected embedding $\bar{{\mathbf{x}}}^{l}=\sum_{k=1}^{K}{\mathbf{c}}_{k}\cdot\,{\mathbf{x}}^{l}_{\varphi}({\mathbf{z}}_{t(i)})[k]\in\mathbb{R}^{d},$ and then quantize it using LFQ (2(Yu et al., 2024): ${\mathbf{x}}^{l}={\mathcal{Q}}_{\mathrm{lfq}}(\bar{{\mathbf{x}}}^{l})\in\{-1,+1\}^{d}.$ We then apply the straight-through estimator (Van Den Oord et al., 2017) to obtain an image $\hat{{\mathbf{x}}}={\mathcal{D}}(\tilde{{\mathbf{x}}}_{\varphi}({\mathbf{z}}_{t})),~\tilde{{\mathbf{x}}}_{\varphi}({\mathbf{z}}_{t})=\bar{{\mathbf{x}}}+\big[{\mathbf{x}}-\bar{{\mathbf{x}}}\big]_{\mathrm{sg}},$ where $\mathrm{sg}$ denotes the stop-gradient operator. Finally, we compute the differentiable likelihood as $q({\mathbf{y}}|\tilde{{\mathbf{x}}}_{\varphi}({\mathbf{z}}_{t}))\propto\exp(-\tfrac{\|{\mathbf{y}}-{\mathcal{A}}(\hat{{\mathbf{x}}})\|_{2}^{2}}{2\sigma^{2}}).$
Discussion. We minimize (3.1) at each time step $t$ to obtain the optimal logits $\varphi^{*}_{t}$. This ensures that the optimized probabilities ${\mathbf{x}}_{\varphi^{*}}({\mathbf{z}}_{t})$ remain close to the prior predictions ${\mathbf{x}}_{\theta}({\mathbf{z}}_{t})$ while aligning with measurements $\mathbf{y}$}. To our knowledge, this is the first use of quantized expectation of codebook embeddings for posterior sampling in discrete diffusion, allowing measurement gradients to backpropagate through $\bar{{\mathbf{x}}}$ and update all entries of ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t})$. This reduces the cost of the measurement term from $\mathcal{O}(K^{L})$ to $\mathcal{O}(KL)$ because (3.1) requires only one forward pass through ${\mathcal{A}}\circ{\mathcal{D}}$ using $\tilde{{\mathbf{x}}}_{\varphi}({\mathbf{z}}_{t})$. We defer complexity analysis to Appendix C.5.
@@ 3.1 Quantized Expectation
3.2 Anchored Remasking
In masked diffusion, the reverse process progressively unmasks tokens. Typical samplers (Austin et al., 2021; Chang et al., 2022; Lou et al., 2024; Sahoo et al., 2024; Shi et al., 2024) choose to unmask based on confidence or random remasking. ADLM (Rout et al., 2025a) shows that prioritizing “anchor” tokens (e.g., nouns or verbs in language, rather than articles or conjunctions) reduces conditional entropy of the remaining tokens in the sequence that subsequently improves generation. To decode anchor tokens, ADLM jointly trains an anchor network in addition to the standard denoising network using an anchored NELBO objective. Here, we propose a training-free variant of anchored denoising, enabling posterior sampling with discrete diffusion.
Let $\varphi$ be a minimizer of (3.1), resulting in ${\mathbf{x}}^{*}_{\varphi}({\mathbf{z}}_{t})=\{({\mathbf{x}}^{*}_{\varphi}({\mathbf{z}}_{t}))^{l}\}_{l=1}^{L}$ that represents categorical distributions over tokens at each position $l\in\{1,\dots,L\}$ at time step $t$. Anchored remasking selects a subset of positions ${\mathcal{P}}_{t}\subseteq\{1,\dots,L\}$ to decode early. The selection is based on the confidence of quantized tokens $\mathbf{x}$} (as defined in §3.1) under the posterior estimate ${\mathbf{x}}^{*}_{\varphi}({\mathbf{z}}_{t})$. Importantly, the posterior estimate is a function of the $L$-length sequence ${\mathbf{z}}_{t}$, and hence encodes the joint relation across all tokens; thus, anchored remasking is a function of all tokens, unlike standard remasking (Chang et al., 2022; Yang et al., 2025b) that depend only on per-token confidence. Formally, we compute the confidence of ${\mathbf{x}}^{l}$ as $\kappa^{l}_{t}=\langle({\mathbf{x}}^{*}_{\varphi}({\mathbf{z}}_{t}))^{l},{\mathbf{x}}^{l}\rangle,$ and choose anchor positions as ${\mathcal{P}}_{t}=\{\,l:\kappa^{l}_{t}\geq\tau_{t}\,\},$ where $\tau_{t}$ is an adaptive threshold that follows from the cosine schedule of MMaDA. We then unmask anchor tokens as: ${\mathbf{z}}_{s}^{l}={\mathbf{x}}^{l}$ with probability $\frac{\alpha_{s}-\alpha_{t}}{1-\alpha_{t}}$ if $l\in{\mathcal{P}}_{t}$ else $\mathbf{m}$}. Once a token is unmasked, it stays unmasked during inference.
Discussion. Diffusion language models tend to be highly confident on low-informative tokens such as articles (“a”, “an” or “the”) or conjunctions (Rout et al., 2025a); similarly, discrete image samplers using independent per-token confidence often unmask background pixels first. In contrast, our method leverages the joint posterior ${\mathbf{x}}_{\varphi}({\mathbf{z}}_{t})$ to identify anchor tokens consistent with the measurements. This leads to earlier decoding of informative tokens (e.g., a bird against a flat background), while aligning with the joint distribution; refer to §C.3 for details. While multi-token sampling from the joint is typically intractable for frameworks like MDLM (Sahoo et al., 2024) or MD4 (Shi et al., 2024), anchored diffusion (Rout et al., 2025a) circumvents this limitation by reducing the conditional entropy of the remaining tokens. We follow a similar anchoring strategy for adaptive decoding of discrete image tokens.
Table 2: Quantitative results on general inverse problems. We report results on two additional linear (random inpainting, motion deblurring) and nonlinear (HDR, nonlinear blur) tasks. Since G2D2 and SGDD do not evaluate on these tasks, we compare our discrete sampler against representative continuous baselines: DPS (pixel-space) (Chung et al., 2023) and PSLD (latent-space) (Rout et al., 2023).
@@ 3.2 Anchored Remasking
4 Experiments
Baselines. Since our focus is on discrete diffusion, we first compare with existing discrete methods: G2D2 (Murata et al., 2024) and SGDD (Chu et al., 2025). To provide a comprehensive evaluation, we also include established continuous baselines, both in pixel space (DPS (Chung et al., 2023), DDRM (Kawar et al., 2022), and DiffPIR (Zhu et al., 2023)) and in latent space (PSLD (Rout et al., 2023) and ReSample (Song et al., 2024)). Thus, our evaluation spans both discrete and continuous paradigms. For language modeling, we compare with LLaDA-8B-Instruct (Nie et al., 2025b) and Dream-7B-Instruct (Ye et al., 2025).
Benchmarks. We evaluate on the standard inverse problem benchmarks used in prior works. For high-resolution faces we use FFHQ at $256\times 256$ (Karras et al., 2019), and for diverse natural images we use ImageNet at $256\times 256$ (Deng et al., 2009). Performance is measured using three standard metrics: Learned Perceptual Image Patch Similarity (Zhang et al., 2018) (LPIPS), Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM) (Wang et al., 2004). All methods are evaluated at the same resolution and on the same images following G2D2 (Murata et al., 2024). For question answering, we use Qwen2.5-3B model (Yang et al., 2025a) to create a benchmark of 20 questions (§ C.8), where evaluation is done using Qwen3-0.6B as a reward model and Qwen3-8B as an independent judge.
[…]
Figure 4: Qualitative results on stylization. We present four style–content combinations. For each case, our APS algorithm conditions on a single reference style image together with a text prompt to generate the stylized output images.
@@ 4.1 Results on Linear Inverse Problems
Evaluation on FFHQ: Table 1 (a) shows that APS outperforms prior discrete diffusion samplers across both super resolution and Gaussian deblurring tasks. For super resolution, APS reduces LPIPS by $13.65\%$ and improves PSNR by $2.11\%$. On Gaussian deblurring, APS lowers LPIPS by $3.83\%$ and raises PSNR by $5.88\%$. The large variant, APS-L, pushes performance even further, achieving up to a $31.36\%$ gain over G2D2 in terms of LPIPS. These results show that our quantized expectation (§3.1) and anchored remasking (§3.2) strategies not only outperform discrete diffusion baselines but often surpass strong continuous diffusion methods, such as DiffPIR (which uses pixel-space diffusion) and PSLD (which uses latent-space diffusion).
Figure 2 (top two rows) presents qualitative comparisons for super resolution ($4\times$) and Gaussian deblurring on FFHQ. DPS produces over-smoothed results with blurry facial details, while G2D2 often introduces artifacts and fails to restore a natural facial structure. Our APS sampler yields sharper textures and more faithful reconstructions, recovering details such as hair strands, facial contours, and eyeglass edges with higher perceptual quality. The large variant, APS-L, further enhances structure and realism, generating photo-realistic outputs with finer details and fewer artifacts. These results confirm that APS and APS-L deliver superior qualitative performance, where perceptual quality is crucial.
Evaluation on ImageNet: Table 1 (b) quantifies that our method achieves consistent improvements across both SR and Gaussian deblurring tasks on the ImageNet benchmark. For super resolution, APS reduces LPIPS by $7.16\%$ compared to G2D2 and improves PSNR by $4.74\%$. On Gaussian deblurring, APS improves PSNR by $8.81\%$ while maintaining comparable LPIPS. APS-L improves LPIPS by up to $35.82\%$ compared to G2D2. These results confirm that APS outperforms prior discrete posterior samplers and often surpasses continuous baselines such as DiffPIR and PSLD with superior perceptual quality and image fidelity.
Figure 2 (bottom two rows) shows SR ($4\times$) and Gaussian deblurring results on ImageNet. Notably, DPS produces overly smooth outputs with a loss of fine details, while G2D2 introduces struggles to recover sharp edges. In contrast, APS reconstructs sharper textures (e.g., the fur of the dog and the feathers of the swan) and yields more natural color. The large variant, APS-L, further enhances structural fidelity, recovering finer details in challenging regions such as a frog’s skin texture and a goat’s fur. These examples highlight that our approach achieves superior perceptual quality and faithful structure reconstruction compared to both continuous and discrete diffusion baselines.
@@ 4.2 Results on General Inverse Problems
Table 2 shows that APS generalizes effectively to more challenging linear (random inpainting, motion deblurring) and nonlinear (HDR, nonlinear blur) inverse problems on FFHQ and ImageNet. Unlike existing methods such as G2D2 and SGDD, which were demonstrated on limited tasks, APS achieves strong perceptual quality (lower LPIPS) and reconstruction fidelity (higher PSNR). For instance, on ImageNet motion deblurring APS-L attains $0.318$ LPIPS and $25.19$ PSNR, substantially outperforming continuous baselines DPS and PSLD. Figure 3 shows the qualitative results. In nonlinear tasks such as HDR and nonlinear blur, APS delivers sharper, more consistent reconstructions, closing the gap with continuous diffusion while operating within a purely discrete framework. These results highlight the broader applicability and robustness of our approach compared to existing discrete diffusion samplers.
@@ 4.3 Results on Reference-based Stylization
Table 3: Quantitative results on stylization. APS enables new capabilities such as reference-based stylization in MMaDA.
@@
We compare APS with the discrete diffusion baseline MMaDA (Yang et al., 2025b) and existing continuous diffusion baselines (shaded gray). Table 3 reports quantitative results, while Figure 4 shows qualitative examples. APS consistently improves over the base MMaDA model across all metrics: ImageReward (Xu et al., 2023), CLIP-T (Radford et al., 2021), and DINO (Caron et al., 2021) demonstrating that our posterior sampler successfully unlocks zero-shot stylization capabilities that are otherwise absent in the base model. By formulating style alignment as a reward maximization problem, APS guides the discrete generation trajectory to capture style attributes from the reference image without requiring auxiliary control networks or fine-tuning. Notably, APS surpasses established continuous baselines such as IP-Adapter (Ye et al., 2023) and StyleAligned (Hertz et al., 2023) on ImageReward and CLIP-T metrics, despite relying on a weaker generative prior.
4.4 Results on Question Answering
Although this paper primarily focuses on discrete diffusion for vision tasks, we show that our approach generalizes to language by evaluating it on a question answering task.
Figure 5 qualitatively demonstrates the efficacy of Anchored Posterior Sampling (APS) in diffusion-based question answering. While Dream-7B-Instruct suffers from incoherence due to a known artifact of independent sampling from the product of marginals, APS employs a lightweight reward model (Qwen3-0.6B) to enable joint sampling. Thus, APS yields structurally consistent and semantically richer outputs, such as better rhyming in the poem.
These qualitative gains are further confirmed by our quantitative evaluation using a held-out oracle judge (Qwen3-8B) as given in Table 4. APS achieves a 21.99% relative improvement over the state-of-the-art Dream-7B-Instruct baseline without additional training. Implementation details are provided in Appendix C.1.
Figure 5: Qualitative comparison on a creative writing prompt. APS produces better rhyming and more structurally coherent creative writing compared to the generic outputs of the baseline.
Table 4: Quantitative results on question answering.
Method Reward Model Oracle Judge
(Qwen3-0.6B) (Qwen3-8B)
LLaDA -0.0970 2.1699
Dream 0.9970 3.7523
Dream+APS (Ours) 1.3786 4.5775
The appendix provides extensive details on our Anchored Posterior Sampling method. We present the theoretical derivation of variational bounds (§B), all implementation specifics and hyperparameters (Algorithm 1, Table 5), and a detailed ablation study (§C.3, Figure 6) demonstrating the impact of our innovations. We also showcase the superior computational efficiency (Table 8) and numerous additional qualitative results (§C) across complex inverse problems, high-resolution text-guided block inpainting (Figure 12), and extended stylization results (Figure 13 and  14).